# Data-driven $h^{2}$ model reduction for linear discrete-time systems

### v2

@@ VI Application to Computational Neuroscience
Therefore, in this experiment, as in [31], we model the relationship between the neural activity of C. elegans, regarded as the state, and changes in oxygen concentration, regarded as the control input, using a state-space model. Furthermore, we demonstrate that the ROMs obtained by the proposed data-driven method can reconstruct the original system more accurately in the sense of the $h^{2}$-norm compared to DMDc [28, 27].
@@ VI-A Problem setting
We consider the following system with noise $\epsilon_{k}$:
[…]
where $k=1,\ldots,L-1$.
First, we describe the measurement data for the original system (34). The data utilized in this study are accessible at https://osf.io/a64uz/ and are linked to two prior experimental papers [43, 31]. Using the data corresponding to [43], we constructed the state data $X$ and $X_{\mathrm{p}}$ with noise and $\tilde{X}$ and $\tilde{X}_{\mathrm{p}}$ without noise, as well as the control input data $U$, as defined by (17) and (34). More specifically, we utilized the calcium imaging data from the npr-1 mutant strain, which is sensitive to changes in oxygen concentration, during the pre-lethargus stage of this strain as the state data. The changes in oxygen concentration were used as the control input. Note that, in this dataset, the parameters for (17) were set as $N=1$, $L=4045$, $n=114$, and $m=1$. Using the MATLAB randn command, the noise $\epsilon_{k}$ was generated. Before generating noise, we fixed the seed with rng(0,’twister’) to ensure exact reproducibility.
@@
[…]
where $x_{k}^{(h_{x})}$, $\tilde{x}_{k}^{(h_{x})}$, and $u_{k}^{(h_{u})}$ represent the $k$-th columns of the matrices $X^{(h_{x})}$, $\tilde{X}^{(h_{x})}$, and $U^{(h_{u})}$, respectively, and $\epsilon_{k}^{(h_{x})}=(\epsilon_{k},\dots,\epsilon_{k+h_{x}-1})$. Using (35), $A^{(h_{x})}$ and $B^{(h_{u})}$ matrices are calculated from $X^{(h_{x})}$, $X^{(h_{x})}_{\mathrm{p}}$, and $U^{(h_{u})}$ as follows:
@@
We then describe other parameters. The initial step-size $\alpha$, Armijo parameter $c_{1}$, and search control parameter $\rho$ required for the back tracking method were $\alpha=1$, $c=10^{-4}$, and $\rho=0.5$, respectively. The tolerance $tol$ was $tol=10^{-3}$.
In this experiment, we demonstrate that the ROMs obtained by the proposed method approximates the true system (35) more accurately in terms of frequency response compared to the ROMs obtained by DMDc. Specifically, the objective function is the relative error $\frac{\|H-\hat{H}\|_{h^{2}}}{\|H\|_{h^{2}}}$ with respect to the transfer functions $H$ and $\hat{H}$ of the original system (35) and ROM, respectively. It should be noted that, while data-driven model reduction methods such as the Loewner framework and Data-driven Balanced Truncation (BT) are well-known, these methods require frequency response data or impulse response data as input. Therefore, they are challenging to apply directly in this context.
@@ VI-B Numerical results
We present the results of applying Algorithm1 to the given state and input data. The initial point for Algorithm 1 is set to the reduced matrices obtained via DMDc, and $X^{(h_{x})},U^{(h_{u})},X^{(h_{x})}_{\mathrm{p}},\hat{A}$ satisfy Assumptions (A1)-(A6).
Figure 1 shows the $h^{2}$-norm plotted for different values of $r$ when the noise coefficient $\beta$ is 00. Similar to conventional model reduction methods, it is confirmed that larger values of $r$ lead to ROMs that better approximate the original system (35). Nevertheless, even at $r=400$ the relative $h^{2}$-error plateaus at about $15\%$ because an intrinsically nonlinear model is being approximated by a purely linear ROM.
Figure 2 illustrates the $h^{2}$-norm for different noise coefficients $\beta$ when $r=200$. When data satisfying the assumptions is used, a better ROM in terms of the $h^{2}$-norm is obtained compared to the initial point provided by DMDc. In the case of higher noise ($\beta=0.01$), the noise impact cannot be ignored, causing the gradient to deviate significantly from the true gradients, which results in limited improvement. Conversely, when the noise level is small ($\beta=0.001$), we observe a 50–60% improvement, similar to the no-noise case ($\beta=0$).
Figure 1: Convergence behavior of (10) with $r$
Figure 2: Convergence behavior of (10) with $\beta$
@@ VII Concluding remarks
Several future directions are outlined below. First, we aim to construct more computationally efficient algorithms for large-scale problems. Second, we plan to report detailed results for additional application domains. Preliminary experiments indicate that the proposed method also outperforms DMDc when modeling daily new infection counts in epidemiology [33]. Third, we will extend (10) to an optimization problem on a Riemannian manifold. As shown in [12] and [44], the reduced matrix $\hat{A}$ obtained at each iteration is inherently stable, so the explicit stability constraint becomes unnecessary.
@@ Acknowledgment
This work was supported by JSPS KAKENHI under Grant Numbers 23K03899 and 25KJ0986.
@@
Since Assumption (A5), $X^{\dagger}X=I_{n}$. From [39], under Assumptions (A1) and (A3), (27) has a unique solution $R_{\text{data}}=R$. Similarly, under Assumptions (A2) and (A3), (28) has a unique solution $S_{\text{data}}=S$. $\Box$
@@
Since (21) and (22) are solutions of (20), $\operatorname{rank}\bigl{(}\begin{smallmatrix}X&U\end{smallmatrix}\bigl{)}=\operatorname{rank}\bigl{(}\begin{smallmatrix}X&U&X_{\mathrm{p}}X^{\top}\end{smallmatrix}\bigl{)}$ holds. Thus, Assumption (A4) implies that (20) has the unique solution $(Z_{A,\text{data}},Z_{B,\text{data}})=(Z_{A},Z_{B})$. $\Box$
@@
(23) has the unique solution (24) when Assumption (A5) holds. From (32), we obtain $US_{B}=U_{B}S_{\text{data}}=UB^{\top}S_{\text{data}}$. Therefore, from Assumption (A6), $U^{\dagger}U=I_{m}$ holds. Thus, (33) holds. $\Box$
@@
Since $Z_{A,\text{data}}=Z_{A}$, $Z_{B,\text{data}}=Z_{B}$, and $U_{B,\text{data}}=U_{B}$ hold from Lemmas 2 and 3, the unique solutions $R$ and $S$ of (8) and (9), respectively, coincide with the unique solutions $R_{\text{data}}$ and $S_{\text{data}}$ of (27) and (28), as stated in Lemma 1.
Furthermore, from Lemma 3 and (8), we have $S^{\top}AR=S^{\top}(R-B\hat{B}^{\top})(\hat{A}^{-1})^{\top}=(S^{\top}R-S_{B}^{\top}\hat{B}^{\top})(\hat{A}^{-1})^{\top}$.
Therefore, under Assumptions (A1)-(A6), the gradients of Proposition 2 are rewritten as (29)-(31). $\Box$
@@
To solve (20) for $(Z_{A},Z_{B})$, it is necessary to solve $NL-N$ linear equations of $(NL-N)\times(n+m)$. Considering the least-squares problem, the complexity is $\mathcal{O}((NL-N)^{2}(n+m))$. Similarly, the complexity required to solve (23) for $U_{B}$ is $\mathcal{O}((NL-N)^{2}n)$. The discrete‐time Lyapunov equations (6) and (7) are each solvable in $\mathcal{O}(r^{3})$ via the Bartels-Stewart method [39]. For the Sylvester equations (27) and (28), forming $X^{\dagger}$ and the matrix products $X^{\dagger}Z_{A}$, $X^{\dagger}Z_{B}\hat{B}^{\top}$, and $X^{\dagger}(X_{\mathrm{p}}-U_{B})$ each require $\mathcal{O}((NL-N)n^{2})$. With $X^{\dagger}X=I_{n}$, the resulting $n\times n$ Sylvester equations are solved in $\mathcal{O}(n^{3})$ using the Bartels-Stewart method [39], giving a total cost $\mathcal{O}((NL-N)n^{2}+n^{3})$. $\Box$

### v3

@@ VI Application to Computational Neuroscience
As in [26], we model the relationship between the neural activity of C. elegans, regarded as the state, and changes in oxygen concentration, regarded as the control input, using the following system with noise $\epsilon_{k}$:
$\begin{aligned} \tilde{x}_{k+1}&=A\tilde{x}_{k}+Bu_{k},\\ x_{k}&=\tilde{x}_{k}+\beta\epsilon_{k},\end{aligned}$ (34)
where $k=1,\ldots,L-1$, and the noise $\epsilon_{k}$ was generated using the MATLAB randn command. Furthermore, we demonstrate that the ROMs obtained by the proposed data-driven method can reconstruct the original system more accurately in the sense of the $h^{2}$-norm compared to DMDc [22, 21].
In experiments, we fixed the seed with rng(0,’twister’) to ensure exact reproducibility. Furthermore, the initial step-size $\alpha_{\text{init}}$, Armijo parameter $c_{1}$, search control parameter $\rho$ required for the backtracking method, and tolerance toltol in Algorithm 1 were $\alpha=1$, $c=10^{-4}$, $\rho=0.5$, and $tol=10^{-3}$ respectively.
@@ VI-A Problem setting
First, we describe the measurement data for the original system (34). The data utilized in this study are accessible at https://osf.io/a64uz/ and are linked to two prior experimental papers [39, 26]. Using the data corresponding to [39], we constructed the state data $X$ and $X_{\mathrm{p}}$ with noise and $\tilde{X}$ and $\tilde{X}_{\mathrm{p}}$ without noise, as well as the control input data $U$, as defined by (18) and (34). More specifically, we utilized the calcium imaging data from the npr-1 mutant strain, which is sensitive to changes in oxygen concentration, during the pre-lethargus stage of this strain as the state data. The changes in oxygen concentration were used as the control input. Note that, in this dataset, the parameters for (18) were set as $N=1$, $L=4045$, $n=114$, and $m=1$.
@@
$\begin{aligned} \tilde{x}_{k+1}^{(h_{x})}&=A^{(h_{x})}\tilde{x}_{k}^{(h_{x})}+B^{(h_{u})}u_{k}^{(h_{u})},\\ x_{k}^{(h_{x})}&=\tilde{x}_{k}^{(h_{x})}+\beta\epsilon_{k}^{(h_{x})}\end{aligned}$ (35)
where $x_{k}^{(h_{x})}$, $\tilde{x}_{k}^{(h_{x})}$, and $u_{k}^{(h_{u})}$ represent the $k$th columns of the matrices $X^{(h_{x})}$, $\tilde{X}^{(h_{x})}$, and $U^{(h_{u})}$, respectively, and $\epsilon_{k}^{(h_{x})}=(\epsilon_{k},\dots,\epsilon_{k+h_{x}-1})$. Using (35), $A^{(h_{x})}$ and $B^{(h_{u})}$ matrices are calculated from $X^{(h_{x})}$, $X^{(h_{x})}_{\mathrm{p}}$, and $U^{(h_{u})}$ as follows:
@@
To assess the effectiveness of the proposed method on the original system (35), we report the relative error $\|H-\hat{H}\|_{h^{2}}/\|H\|_{h^{2}}$ where $H$ denotes the transfer functions of the original system (35). It should be noted that, data-driven model reduction methods (e.g., data-driven BT) that use frequency response data or impulse response data as input are difficult to apply directly in this context.
@@ VI-B Numerical results
We present the results of applying Algorithm 1 to the given state and input data. The initial point for Algorithm 1 is set to the reduced matrices obtained via DMDc, and $X^{(h_{x})},U^{(h_{u})},X^{(h_{x})}_{\mathrm{p}},\hat{A}$ satisfy (A1)-(A6).
Fig. 1 (upper) shows the $h^{2}$-norm plotted for different values of $r$ when the noise coefficient $\beta$ is 0. Similar to conventional model reduction methods, it is confirmed that larger values of $r$ lead to ROMs that better approximate the original system (35). Nevertheless, even at $r=400$ the relative $h^{2}$-error plateaus at about $15\%$ because an intrinsically nonlinear model is being approximated by a purely linear ROM.
Fig. 1 (lower) illustrates the $h^{2}$-norm for different noise coefficients $\beta$ when $r=200$. When data satisfying the assumptions is used, a better ROM in terms of the $h^{2}$-norm is obtained compared to the initial point provided by DMDc. In the case of higher noise ($\beta=0.01$), the noise impact cannot be ignored, causing the gradients to deviate significantly from the true gradients, which results in limited improvement. Conversely, when the noise level is small ($\beta=0.001$), we observe a 50–60% improvement, similar to the no-noise case ($\beta=0$).
Figure 1: Convergence behavior of (11): (upper) variation with $r$ ($\beta=0$, i.e., no noise), (lower) variation with $\beta$ ($r=200$).
@@ VII Concluding remarks
Future work includes: (i) more computationally efficient algorithms for large-scale problems; (ii) detailed evaluations in additional domains—preliminary results show our method also outperforms DMDc for daily infection modeling in epidemiology [28]; and (iii) a Riemannian reformulation of (11), where $\hat{A}$ is inherently stable and the explicit stability constraint becomes unnecessary [12, 40].
@@ Acknowledgment
This work was supported by JSPS KAKENHI under Grant Numbers 23K28369 and 25KJ0986.
@@
Since (A5), $X^{\dagger}X=I_{n}$. From [36], under (A1) and (A3), (27) has a unique solution $R_{\text{data}}=R$. Similarly, under (A2) and (A3), (28) has a unique solution $S_{\text{data}}=S$. $\Box$
@@
Since (22) are solutions of (21), $\operatorname{rank}\!\begin{pmatrix}X&U\end{pmatrix}=\operatorname{rank}\!\begin{pmatrix}X&U&X_{\mathrm{p}}X^{\top}\end{pmatrix}$ holds. Thus, (A4) implies that (21) has the unique solution $(Z_{A,\text{data}},Z_{B,\text{data}})=(Z_{A},Z_{B})$. $\Box$
@@
(23) has the unique solution (24) when (A5) holds. From (32), we obtain $US_{B}=U_{B}S_{\text{data}}=UB^{\top}S_{\text{data}}$. Therefore, from (A6), $U^{\dagger}U=I_{m}$ holds. Thus, (33) holds. $\Box$
@@
Since $Z_{A,\text{data}}=Z_{A}$, $Z_{B,\text{data}}=Z_{B}$, and $U_{B,\text{data}}=U_{B}$ hold from Lemmas 2 and 3, the unique solutions $R$ and $S$ of (9) and (10), respectively, coincide with the unique solutions $R_{\text{data}}$ and $S_{\text{data}}$ of (27) and (28), as stated in Lemma 1. Furthermore, from Lemma 3 and (9), we have $S^{\top}AR=S^{\top}(R-B\hat{B}^{\top})(\hat{A}^{-1})^{\top}=(S^{\top}R-S_{B}^{\top}\hat{B}^{\top})(\hat{A}^{-1})^{\top}$. Therefore, under (A1)-(A6), the gradients of Proposition 2 are rewritten as (29)-(31). $\Box$
Proof of Proposition 3.
The sets ${\mathcal{E}}_{\mathrm{s}}$ by (4) (open; e.g. [25]) and ${\mathcal{E}}_{\mathrm{n}}:=\{\hat{\theta}\mid\det\hat{A}\neq 0\}$ (open by continuity of $\det$) are open, hence ${\mathcal{E}}=\mathcal{E}_{\mathrm{s}}\cap\mathcal{E}_{\mathrm{n}}$ is open. Thus, there is $\delta>0$ with $\{\bar{\theta}\mid\|\bar{\theta}-\hat{\theta}_{\ell}\|_{F}<\delta\}\subset\mathcal{E}$. Therefore, when $\tilde{\nabla}f(\hat{\theta}_{\ell})\neq 0$, if $\alpha^{\star}_{\ell}:=\varepsilon/\|\tilde{\nabla}_{\hat{A}}f(\hat{A}_{(\ell)})\|_{F}$, then for any $\alpha\in(0,\alpha^{\star}_{\ell})$, $\bar{\theta}\in\mathcal{E}$ holds. Note that when $\tilde{\nabla}f(\hat{\theta}_{\ell})=0$, $\bar{\theta}=\hat{\theta}$. $\Box$
@@
[…]
We define the closure $K:=\overline{\{\hat{\theta}_{\ell}\mid\ \ell\geq 1\}}$ of the sequence generated by Algorithm 1. Since $\{\hat{\theta}_{\ell}\}$ is bounded and the limit $\lim_{\ell\to\infty}\hat{A}_{(\ell)}$ satisfies (A3), $K$ is a compact subset of $\mathcal{E}$}. Furthermore, since $f$ is smooth, $\nabla^{2}f$ is continuous; hence, by the extreme value theorem, $L_{\nabla}:=\sup_{x\in K}\|\nabla^{2}f(x)\|<\infty$. Therefore,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\nabla f(x)-\nabla f(y)\|\leq L_{\nabla}\|x-y\|\quad\forall\,x,y\in K.}$ (36)
[…]
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f(\hat{\theta}_{\ell+1})\leq f(\hat{\theta}_{\ell})-\frac{c_{1}}{\alpha_{\ell}}\|s_{\ell}\|^{2}\leq f(\hat{\theta}_{\ell})-\frac{c_{1}}{\alpha_{\text{init}}}\|s_{\ell}\|^{2}.}$ (37)
[…]
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\nabla f(\hat{\theta}_{\ell+1})\|\leq\|\nabla f(\hat{\theta}_{\ell})\|+L_{\nabla}\|s_{\ell}\|\leq(\underline{\alpha}^{-1}+L_{\nabla})\|s_{\ell}\|,}$ (38)
where the first inequality follows from (36), and the second from the definition of $s_{\ell}$.
Lastly, $\{\hat{\theta}_{\ell}\}$ is bounded and, by the Bolzano–Weierstrass theorem, there exists a convergent subsequence $\{\hat{\theta}_{\ell_{j}}\}$ whose limit lies in $K$. Furthermore, since $f$ is real-analytic, it is a KL function. Hence, by (37), (38), and [37, Thm. 3.2], Algorithm 1 converges to a stationary point in $K$. $\Box$
Proof of Theorem 3.
To solve (21) for $(Z_{A},Z_{B})$, it is necessary to solve NLNNL-N linear equations of $(NL-N)\times(n+m)$. Considering the least-squares problem, the complexity is $\mathcal{O}((NL-N)^{2}(n+m))$. Similarly, the complexity required to solve (23) for $U_{B}$ is $\mathcal{O}((NL-N)^{2}n)$. The discrete‐time Lyapunov equations (7) and (8) are each solvable in $\mathcal{O}(r^{3})$ via the Bartels-Stewart method [36]. For the Sylvester equations (27) and (28), forming $X^{\dagger}$ and the matrix products $X^{\dagger}Z_{A}$, $X^{\dagger}Z_{B}\hat{B}^{\top}$, and $X^{\dagger}(X_{\mathrm{p}}-U_{B})$ each require $\mathcal{O}((NL-N)n^{2})$. With $X^{\dagger}X=I_{n}$, the resulting $n\times n$ Sylvester equations are solved in $\mathcal{O}(n^{3})$ via the Bartels-Stewart method [36], giving a total cost $\mathcal{O}((NL-N)n^{2}+n^{3})$. $\Box$